Two objects of mass M ar connected by a light rigid rod of length L.The system is placed on a horizontal surface.An impulse J=MV is applied to one of the objects perpendicular to the rod.Calculate the angular acceleration of the system just after impact.(NEGLECT FRICTION)

Grade:12th pass

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Dear Student,

Angular momentum impulse given = MV(L/2) = Isystem w
MV(L/2) = (2ML2/4) w
Angular velocity just after the impact = w = 2V/L
….….….….….….….….….….….….….….….….….….….….…....
Thanks